# Optical-Cavity Manipulation Strategies of Conical Intersections Mediated Singlet Fission Systems

### v1

@@ Abstract
$\rm Molecule$
$\Delta\rm G(meV)$
$\epsilon_{\rm TT}(\%)$
$\tilde{\epsilon}_{\rm TT}/\epsilon_{\rm TT}$
Ref.
$\rm Tetracene$
$150$
0.01
$\sim 16$
Mart
$\rm Pentacene$
$-110$
100
$\sim 1.4$
Mart
$\rm Hexacene$
$-630$
48
$\sim 3.3$
Mart
$\rm Rubrene$
$50$
17,40,>99.917,40,>99.917 , 40 , > 99.9
$\sim 1.3$
Takahashi ; Maslennikov ; Malin13
@@
In this Perspective, we summarize the status quo of the studies into the CI-driven cavity-mediated SF from the point of view of theoreticians. We begin with the overview of the fundamental processes shaping the SF mechanisms and outline spectroscopic signatures of these mechanisms. Then we put these findings in the general context of theoretical studies of PPP molecular systems and materials, outlining the open questions and computational challenges. For making the present work reader-friendly, we choose to concentrate on a single representative SF system, viz. rubrene dimers (RDs). All simulations of the present work were performed with the multiple Davydov Ansatz (mDA) method zhao2 ; zhao1 which is a well-established many-body wave-function-based variational integrator of quantum dynamics employing Gaussian coherent states (see the last section of this work for the discussion of the suitable numerical tools for the simulation of PPP dynamics and spectroscopy). If the mDA multiplicity is high enough (this is so for all numerical illustrations of the present work), the mDA can be regarded as the numerically accurate method which faithfully reproduces evolutions of quantum systems with many DOFs.
@@
Polaritonic CI landscapes. In rubrene, the energy of the TT state, $\rm E_{TT}$ is slightly higher than the energy of the singlet state, $\rm E_{S_{1}}$. Hence SF – without thermal activation Miyata or without excitation to higher-lying electronic states $\rm E_{S_{n}}$ Malin12 – is rather inefficient. Furthermore, the usual mechanism of polaritonic enhancement of SF through the cavity-finetuned resonance of the LP and TT states, is ineffective in rubrene, too Ribeiroa ; Takahashi ; Mart . Hence rubrene is a suitable system for exploring novel cavity-induced mechanisms of the SF enhancement.
@@
Figure 2: Cuts of the adiabatic PESs along the tuning mode ($Q^{(j)}_{c}=0$) for one (left column; panels a, b) and two (right column; panels c, d) RDs in the cavity for the first few polaritonic manifolds $N$. (a) No coupling to the cavity mode ($\Omega=0$), the bare CI is located in the Franck-Condon region at $Q^{(1)}_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$. (b, c, d) Strong coupling to the cavity mode ($\Omega=0.2\leavevmode\nobreak\ \rm eV$). The fraction of the TT states ($W_{\rm TT}$ of Eq. (3)) is indicated through the line color coding of PESs (a, c), while fraction of the photonic states ($W_{\rm ph}$ of Eq. (2)), is indicated through the line color coding of the PESs (b, d). The dash lines in panel (b) show positions of the polaritonic CIs.
Following Ref. sun1 , we consider a microcavity with one or two RDs placed into it. To focus on the net effect of the cavity-induced control mechanisms, the loss of photons in the cavity is neglected. This is justified, since the ultrahigh-Q PhCnB cavities ($\rm Q\sim 750,000$) have been created by using a five-hole taper design Deotare and the timescale of the studied processes (hundreds of femtoseconds) is quite short.
@@
In this section, the $j$th RD participating in the CI-mediated SF is modeled as a system with three electronic states (the ground state ${\rm S}_{0}^{(j)}$, the singlet excited state ${\rm S}_{1}^{(j)}$, and the correlated triplet pair state ${\rm TT}^{(j)}$) and two vibrational modes (the tuning mode with a frequency $\Omega_{t}=0.186$ eV and coordinate $Q^{(j)}_{t}$, and the coupling mode with a frequency $\Omega_{c}=0.0154$ eV and coordinate $Q^{(j)}_{c}$iyata ; sun1 . The electronic energies of the singlet and triplet states are fixed at $E_{{\rm S}_{1}}^{(j)}=2.23$ eV and $E_{{\rm TT}_{1}}^{(j)}=2.28$ eV. The cavity is modelled via a single photonic mode of a frequency $\omega_{c}=2.256$ eV, which is nearly in the electronic resonance with the ${\rm S}_{1}^{(j)}$ and ${\rm TT}_{1}^{(j)}$ states. $\Omega$ determines the coupling of the RDs to the cavity mode. If not explicitly stated otherwise, we set $\Omega=0.2$ eV, which corresponds to the strong cavity-RD coupling regime. To increase space of the adjustable parameters, the rotating-wave approximation (RWA) in the description of the cavity-matter interaction was not applied.
@@
where $|\Psi_{n}({\bf Q})\rangle$ is the adiabatic eigenfunction of the polaritonic Hamiltonian. $N$ specifies the excitation manifold of the polaritonic PESs. For example, $N=0$ corresponds to the collective ground state, $N=1$ yields the singly-excited manifold, etc. To characterize and visualize various states, it is elucidating to introduce several quantifiers. For instance,
@@
The fractions of the TT states ($W_{\rm TT}$ of Eq. (3)) and photonic states ($W_{\rm ph}$ of Eq. (2)) give a convenient visualization of the polaritonic PESs, which is implemented through the line color coding in Figure 2. Panels (a, b) represent a single RD in the cavity, and show the cut of the polaritonic PESs along the symmetric tuning mode $Q_{t}$ at $Q_{c}=0$. Panel (a) corresponds to a cavity-free RD. Hence, ${\rm S}_{1}^{(1)}$ and ${\rm TT}^{(j)}$ are not dressed with the cavity mode, and only the intrinsic RD CI at $Q_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$ exists in this case. It follows that this CI located in the Franck-Condon region provides the only SF pathway. In the presence of cavity, the PESs change substantially, and cascades of polaritonic CIs appear in Figure 2(b). As the coupling to the cavity increases, the two polaritonic CIs move away from the Franck-Condon point, the distance between the pair of polaritonic CIs increases with $N$, and positions of these CIs lie along the V-shaped dash lines in Figure 2(b). It is essential that the fraction of the hybrid matter-photon PESs increases with $N$ near the Franck-Condon point. This facilitates the cavity-assisted SF.
@@
SF Dynamics Mediated by Polaritonic CIs. Let us explore how the polaritonic PESs coupled through cascades of the cavity-induced CIs affect the SF dynamics. For this to achieve, it is worthwhile to study what is the impact of the initial excitation of the photonic mode on the ensuing singlet-to-triplet population transfer. In our simulations, the cavity mode is assumed to be initially, at $t=0$, in the coherent state
@@
Figure 3: Population dynamics of the triplet state (upper panels) and average photon mode populations $P_{cav}(t)$ (lower panels) for several initial photon-mode pumping numbers $|\mu_{1}(0)|^{2}=N$. Left column (panels a, c): one RD. Right column (panels b, d): two RDs.
Let us begin with the considerations of the upper panels, which show the triplet populations for different $N$. For a single RD, the triplet population – on the average – increases with $N$ up to $N=6$. For example, the population of $\approx 0.28$ is observed at 300 fs for $N=6$, which is approximately twice larger than the population obtained in the same model without a cavity. However, the SF efficiency starts to decrease at $N=7$ and is fully suppressed for $N\geq 8$. This effect can rationalized from the different points of view. On the one hand, it can be viewed as excitation localization on the singlet state ${\rm S}_{1}^{(1)}$ and on the photonic mode, which renders the ${\rm TT}^{(1)}$ effectively decoupled and the SF switched off (see Refs. Localization19 ; Localization20 for engineering of long-lived localized states in the cavity-QED systems). A qualitatively similar localization is realized, for example, in the spin-boson model in the regime of the strong coupling of the spin system to the bosonic bath Wu , even at different system preparations LP23 . Alternatively, the effect can be envisaged as an effective annihilation of the nonadiabatic coupling between the singlet and triplet states Gelin09 . It operates through the combined effect of the cavity field (through Rabi cycling) the system itself (through the intrinsic ${\rm S}_{1}^{(1)}-{\rm TT}^{(1)}$ CI), since the photonic mode does not directly couple the singlet and triplet states. Note also that the population evolution of the triplet states starts not from $t=0$, but from a characteristic time $\tau_{N}$ which increases with $N$. This indicates that the wavepacket needs $\approx\tau_{N}$ to reach the polaritonic CIs.
@@
Let us now inspect evolutions $P_{cav}(t)$ of the cavity-mode populations vs. the initial photon pumping numbers $N$ ($P_{cav}(0)=N$) which deliver a complementary information. $P_{cav}(t)$ for a single RD is depicted in Figure 3(c). For $N$ from 1 to 6, the populations exhibit similar behavior: During the first 50 fs, they show several oscillations with a rough period of 22 fs which is related to the tuning mode. Then they reach a quasi-steady-state $\langle P_{cav}\rangle$ value at longer times. It is worth mentioning that $\langle P_{cav}\rangle$ are slightly larger than the initial populations $P_{cav}(0)=N$. It reveals the population redistribution due to the coupling of the cavity mode and the ${\rm S}_{1}^{(1)}$ state. $P_{cav}(t)$ for $N=7$ shows high-amplitude oscillations during the first $\sim 40$ fs, which matches the time $\tau_{7}$ during which the corresponding triplet population in Figure 3(a) is almost zero. For $N=8$, the high-amplitude oscillations persist over the entire evolution of $P_{cav}(t)$, which is a signature of the localization of the population on the photonic mode and the singlet state, which results in the SF stopping. Interestingly, the period of these fast oscillations reveals the effective Rabi frequency $\Omega_{N}\approx\Omega\sqrt{N}$ ($2\pi/\Omega_{8}\approx 7.3$ fs), which is defined under the assumption that the effective coupling to the cavity scales $\sim\sqrt{N}$.
Figure 3(d) is a counterpart of Figure 3(c), but for two RDs coupled to the cavity. For $N$ from 1 to 6, $P_{cav}(t)$ reach their first local minima at $t\approx 17$ fs (dashed line in panel d), which indicates that a fraction of the photonic-mode population is transferred to the doubly excited electronic states of the RDs. Then, $P_{cav}(t)$ exhibit oscillations with a period of $\sim 26$ fs, which cannot be directly related to the model parameters. In general, excitation to $1\leq N\leq 6$ reflects population transfer between photonic mode and excited states of the RDs. For $7\leq N\leq 9$, $P_{cav}(t)$ shows fast Rabi oscillations at the initial stage, which correlate with the absence of the triplet excitation at the same time interval (cf. Figure 3(b)). For $N=10$, fast Rabi oscillations spread over the entire $P_{cav}(t)$ evolution, which is the signature of the SF switching-off in Figure 3(b). In the long-time limit, the $P_{cav}(t)$ reach quasi-steady states $\langle P_{cav}\rangle$ the values of which are lower than the initial values $P_{cav}(0)$. This indicates a net population transfer from the photonic mode to the doubly-excited electronic states of RDs.
@@
Figure 4: The level structure for two rubrene dimers in the cavity. The superscripts $(1)$ and $(2)$ number the dimers.
@@
The above analyses demonstrate that the SF efficiency is primarily determined by two factors. The first factor is the position of the polaritonic CIs (see Figure 2). If the CIs are located in the Franck-Condon region, the subsequent population transfer is facilitated and the higher SF efficiency is achieved. The second factor is the initial wavepacket preparation and the state of the cavity which is controlled by the photonic-mode pumping number $|\mu_{1}(0)|^{2}=N$. This later parameter plays the crucial role, increasing the SF efficiency for small $N$ and fully suppressing SF once $N$ reaches a certain critical value.
@@
Spectroscopic Signatures of Polaritonic CIs. As demonstrated in the previous sections, strong coupling of RDs to the cavity mode drastically alters the SF efficiency owing to the cohesive action of the cavity-induced and intrinsic CIs. This processes are imprinted into nonlinear femtosecond spectroscopic responses of these systems. To find spectroscopic signatures of the polaritonic SF occurring in multidimensional PES landscapes, the signals must be accurately simulated and adequately interpreted. For this to achieve, it is essential to realistically model RDs. Indeed, all third-order spectroscopic responses consist of three contributions, the ground-state bleach (GSB), the stimulated emission (SE), and the excited-state absorption (ESA) Darius04 ; Leonas . The GSB and SE reflect the photoinduced processes in the manifolds with $N=0$ (the ground state) and $N=1$ (singly-excited states), while ESA involves also the doubly-excited states, $N=2$. Hence modeling individual RDs as systems containing only singly-excited states ${\rm S}_{1}^{(j)}$ and ${\rm TT}^{(j)}$ is insufficient: we have to include higher-excited states ${\rm S}_{n}^{(j)}$ and ${\rm TT}_{n}^{(j)}$ sun2 . Following Miyata ; sun2 , we set $E_{{\rm S}_{n}}^{(j)}=4.33$ eV and $E_{{\rm TT}_{n}}^{(j)}=4.68$ eV in our simulations. For two RDs in the cavity, we thus obtain a single ground state, 5 singly-excited states and 15 doubly-excited states which are specified in Fig. 4. The states, in turn, can be subdivided into the matter states ($n_{c}=0$, left part of Fig. 4) and hybrid states ($n_{c}>0$, right part of Fig. 4).
Here we consider electronic two-dimensional (2D) spectra, because electronic 2D spectroscopy (2DES) is one of the most powerful and information-rich third-order spectroscopic techniques (see Scholes22 and references therein). Qualitatively, 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ can be envisaged as a $T_{w}$-dependent 2D $(\omega_{\tau},\omega_{t})$ maps, which reveal correlations between the states of the system excited at a frequency $\omega_{\tau}$ at time zero and probed at a frequency $\omega_{t}$ at time $T_{w}$. 2DES spectra $S(\omega_{\tau},T_{w},\omega_{t})$ simulated for the two RD system strongly coupled to the cavity mode are presented in Figure 5 for different population times $T_{w}$. From top to bottom, shown are the SE (row (a)), GSB (row (b)), and ESA (row (c)) contributions, as well as the total signal (row (d)).
The SE contribution (row (a)) possesses two pronounced diagonal peaks (DPs) and two cross-peaks (CPs) (similar situation takes place in the purely polaritonic dimer system Bondarev13 ). The DP1 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and DP2 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) reveal the middle polariton (MP) and the UP bands, respectively. The low-intensity LP band cannot be reliably identified in the SE 2DES spectrum, because 2DES peak intensities are proportional to the product of four relevant transition dipole moments, in contrast with absorption spectra, which are proportional to the product of two transition dipole moments (see the discussion in Ref. Jindra15 ). The CP1 ($\omega_{\tau}=2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$) and CP2 ($\omega_{\tau}=2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) quantify the coupling and polaritonic coherence between the states MP and UP. A weak CP ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) at $T_{w}=16$ fs reveals the phonon sideband. The shape and intensity of the DPs and CPs depend significantly on the waiting time $T_{w}$. The DPs at $T_{w}=0$ are predominately elongated perpendicular to the main diagonal, but their elongation turns into parallel at longer population times. A certain similarity of the DPs at $T_{w}=0$ and 32 fs shows a significant contribution of the polaritonic beating with a period of $\sim 32\leavevmode\nobreak\ {\rm fs}$ revealing the energy difference of $0.132\,{\rm eV}$ between the states MP and UP. The two negative CPs are elongated parallel to the main diagonal. Similar to the DPs, the CPs are also significantly modulated by the MP-UP energy-gap beating. The overall decreases of the intensity of the SE contribution with population time mirrors the CI-mediated population transfer from the bright $S_{1}^{(j)}$ states to the dark $TT^{(j)}_{1}$ states and reflects thereby the SF dynamics.
The GSB contribution is shown in panels (b) of Figure 5. It reveals the same pairs of the polaritonic DPs and CPs and, additionally, a pair of weaker polariton-phonon CPs, ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.429\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}=2.6\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}=2.297\leavevmode\nobreak\ {\rm eV}$). In comparison with the SE peaks, their GSB counterparts look rather static. Nevertheless, the GSB contribution reveals the vibrational wave-packet motion in the electronic ground state, which is clearly manifested in the peak evolutions (see Ref. sun2 ).
The ESA contributions are depicted in panels (c) of Figure 5. The SE and ESA contributions to 2DES spectra reveal photophysical processes in the manifold $N=1$ skw20a . However, the two contributions employ different reporter states for delivering the same information: the SE employs the electronic ground state $N=0$, while the ESA employs the doubly-excited manifold $N=2$. Hence shapes and evolutions of predominantly positive SE contributions and predominantly negative ESA contributions are quite similar. Nevertheless, positions of the DPs and CPs in the SE and ESA spectra are slightly shifted with respect to each other. At $T_{w}=0\leavevmode\nobreak\ {\rm fs}$, for example, the (nearly) diagonal ESA peaks are located at ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$) and ($\omega_{\tau}\approx 2.429\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.416\,{\rm eV}$). All ESA peaks are elongated parallel to the $\omega_{t}$ axis, due to a high density of states in the manifold $N=2$. The peak ($\omega_{\tau}\approx 2.297\leavevmode\nobreak\ {\rm eV}$, $\omega_{t}\approx 2.317\leavevmode\nobreak\ {\rm eV}$), which revels transitions from the MP state to the states of the doubly-excited manifold, has the highest intensity. Note also that the relative separation between the DPs and CPs along the $\omega_{t}$ axis changes with $T_{w}$: cf. the spectra at 0 fs with those at 16 fs. This modulation is caused by the superposition of polaritonic and purely vibrational tuning-mode oscillations. This reflects a higher complexity of the ESA contribution, which involves polaritonic states with $N=1$ and 2.
Finally, the total 2DES signal $S(\omega_{\tau},T_{w},\omega_{t})$ is shown in Figure 5(d). Since positions of the DPs and CPs in the SE, GSB and ESA contributions nearly coincide, the total 2DES spectrum exhibits positive, negative and dispersed peak shapes, the amplitudes of which are determined by subtle cancellation effects in the superposition of positive SE and GSB contributions and negative ESA contribution. The overall structure of the total 2DES spectra is dominated by the quartet of polaritonic DPs and CPs. In addition, a number of polariton-vibrational peaks around $\omega_{\tau}\,\mathrm{or}\,\omega_{t}=$ 2.6 eV is seen at different $T_{w}.$ It is remarkable that these hybrid polariton-vibrational peaks are more pronounced in the total signal than in the SE, GSB, and ESA contributions – owing to the aforementioned cancellation effects. The DPs are positive while the CPs exhibit unusual asymmetry: the upper CP is always positive, but the spectral structures around the lower CP can be negative (see the signals for $T_{w}=0$ and 16 fs).

### v2

@@ III CI-mediated SF in optical cavities
$\rm Molecule$ $\Delta\rm G(meV)$ $\epsilon_{\rm TT}(\%)$ $\tilde{\epsilon}_{\rm TT}/\epsilon_{\rm TT}$ Ref.
$\rm Tetracene$ $150$ 0.01 $\sim 16$ Mart
$\rm Pentacene$ $-110$ 100 $\sim 1.4$ Mart
$\rm Hexacene$ $-630$ 48 $\sim 3.3$ Mart
$\rm Rubrene$ $50$ 17,40,>99.917,40,>99.917 , 40 , > 99.9 $\sim 1.3$ Takahashi ; Maslennikov ; Malin13
@@
III.1 Polaritonic CI landscapes
@@
In rubrene, the energy of the TT state, $\rm E_{TT}$, is slightly higher than the energy of the singlet state, $\rm E_{S_{1}}$. Hence SF – without thermal activation Miyata or excitation to higher-lying electronic states $\rm E_{S_{n}}$ Malin12 – is rather inefficient. Furthermore, the usual mechanism of polaritonic enhancement of SF through the cavity-finetuned resonance of the LP and TT states, is ineffective in rubrene, too Ribeiro2018 ; Takahashi ; Mart . Hence rubrene is a suitable system for exploring novel cavity-induced mechanisms of the SF enhancement.
@@
Figure 5: Cuts of the adiabatic PESs along the tuning mode ($Q^{(j)}_{c}=0$) for one (left column; panels a, b) and two (right column; panels c, d) RDs in the cavity for the first few polaritonic manifolds $N$. (a) No coupling to the cavity mode ($\Omega=0$), the bare CI is located in the Franck-Condon region at $Q^{(1)}_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$. (b, c, d) Strong coupling to the cavity mode ($\Omega=0.2\leavevmode\nobreak\ \rm eV$). The fraction of the TT states ($W_{\rm TT}$ of Eq. (18)) is indicated through the line color coding of PESs (a, c), while fraction of the photonic states ($W_{\rm ph}$ of Eq. (17)), is indicated through the line color coding of the PESs (b, d). The dash lines in panel (b) show positions of the polaritonic CIs. Adapted from sun1 . Copyright American Chemical Society.
@@
Following Ref. sun1 , we consider a microcavity with one or two RDs placed into it. To focus on the net effect of the cavity-induced control mechanisms, the loss of photons in the cavity is neglected. This approximation is reasonable. For example, ultrahigh-Q PhCnB cavities with a quality factor of approximately $\rm Q\sim 750,000$ have been fabricated through the utilization of a five-hole taper design Deotare . Moreover, the nanobeam cavity has been optimized to ensure a quality factor of 1700 and a transmission efficiency of $90\%$. When integrated with the bowtie, the hybrid photonic-plasmonic cavity attains a quality factor of 800 and a transmission rate of $20\%$ Conteduca . Thus, the timescale of the studied SF processes (hundreds of femtoseconds) is short in comparison with the lifetimes of such cavities. In this setup, the $j$th RD participating in the CI-mediated SF is modeled as a system with three electronic states (the ground state ${\rm S}_{0}^{(j)}$, the singlet excited state ${\rm S}_{1}^{(j)}$, and the correlated triplet pair state ${\rm TT}^{(j)}$) and two vibrational modes (the tuning mode with a frequency $\Omega_{t}=0.186$ eV and coordinate $Q^{(j)}_{t}$, and the coupling mode with a frequency $\Omega_{c}=0.0154$ eV and coordinate $Q^{(j)}_{c}$iyata ; sun1 . The electronic energies of the singlet and triplet states are fixed at $E_{{\rm S}_{1}}^{(j)}=2.23$ eV and $E_{{\rm TT}_{1}}^{(j)}=2.28$ eV. The cavity is modelled via a single photonic mode of a frequency $\omega_{c}=2.256$ eV, which is nearly in the electronic resonance with the ${\rm S}_{1}^{(j)}$ and ${\rm TT}_{1}^{(j)}$ states. $\Omega$ determines the coupling of the RDs to the cavity mode. If not explicitly stated otherwise, we set $\Omega=0.2$ eV, which corresponds to the strong cavity-RD coupling regime. To increase space of the adjustable parameters, the RWA in the description of the cavity-matter interaction was not applied.
@@
where $|\Psi_{n}({\bf Q})\rangle$ is the adiabatic eigenfunction of the polaritonic Hamiltonian (since the cavity mode is in electronic resonance with the bright RD state, the RWA is almost exact and the number of excitation is nearly conserved). $N$ specifies the excitation manifold of the polaritonic PESs. For example, $N=0$ corresponds to the collective ground state, $N=1$ yields the singly-excited manifold, etc. To characterize and visualize various states, it is elucidating to introduce several quantifiers. For instance,
@@
The fractions of the TT states ($W_{\rm TT}$ of Eq. (18)) and photonic states ($W_{\rm ph}$ of Eq. (17)) give a convenient visualization of the polaritonic PESs, which is implemented through the line color coding in Figure 5. Panels (a, b) represent a single RD in the cavity, and show the cut of the polaritonic PESs along the symmetric tuning mode $Q_{t}$ at $Q_{c}=0$. Panel (a) corresponds to a cavity-free RD. Hence, ${\rm S}_{1}^{(1)}$ and ${\rm TT}^{(j)}$ are not dressed with the cavity mode, and only the intrinsic RD CI at $Q_{t}=0.07,E=2.256\leavevmode\nobreak\ \rm eV$ exists in this case. It follows that this CI located in the Franck-Condon region provides the only SF pathway. In the presence of cavity, the PESs change substantially, and cascades of polaritonic CIs appear in Figure 5(b). As the coupling to the cavity increases, the two polaritonic CIs move away from the Franck-Condon point, the distance between the pair of polaritonic CIs increases with $N$, and positions of these CIs lie along the V-shaped dash lines in Figure 5(b). A very similar situation takes place in the pentacene dimer, in which the polaritonic CI is pushed away from the Franck-Condon region, suppressing the SF process Gu2 . However, it is essential that the fraction of the hybrid matter-photon PESs increases with $N$ near the Franck-Condon point. This reversely facilitates the cavity-assisted SF.
@@
III.2 SF dynamics Mediated by Polaritonic CIs
Let us explore how the polaritonic PESs coupled through cascades of the cavity-induced CIs affect the SF dynamics. Vendrell suggested the number of photons in the cavity, $N$, can be used as a “control knob”, since stimulated emission induced by a cavity resonantly coupled to a specific molecular reaction coordinate strongly depends on $N$ Vendrell1 . To achieve this, it is worthwhile to study the impact of the initial excitation of the photonic mode on the ensuing singlet-to-triplet population transfer. In our simulations, the cavity mode is assumed to be initially, at $t=0$, in the coherent state
@@ III.2 SF dynamics Mediated by Polaritonic CIs
Figure 6: Population dynamics of the triplet state (upper panels) and average photon mode populations $P_{\rm cav}(t)$ (lower panels) for several initial photon-mode pumping numbers $|\mu_{1}(0)|^{2}=N$. Left column (panels a, c): one RD. Right column (panels b, d): two RDs. Adapted from sun1 . Copyright American Chemical Society.
Let us begin with the considerations of the upper panels, which show the triplet populations for different $N$. For a single RD, the triplet population – on the average – increases with $N$ up to $N=6$. For example, the population of $\approx 0.28$ is observed at 300 fs for $N=6$, which is approximately twice larger than the population obtained in the same model without a cavity. As long as these photons remain in the cavity and are available for the next “$\rm S_{0}\to S_{1}\to TT\to charge\leavevmode\nobreak\ transfer$” cycle, the SF efficiency will increase. However, the SF efficiency starts to decrease at $N=7$ and is fully suppressed for $N\geq 8$. This effect can rationalized from the different points of view. On the one hand, it can be viewed as excitation localization on the singlet state ${\rm S}_{1}^{(1)}$ and on the photonic mode, which renders the ${\rm TT}^{(1)}$ effectively decoupled and the SF switched off (see Refs. Localization19 ; Localization20 for engineering of long-lived localized states in the cavity-QED systems). A qualitatively similar localization is realized, for example, in the spin-boson model in the regime of the strong coupling of the spin system to the bosonic bath Wu , even at different system preparations LP23 . Alternatively, the effect can be envisaged as an effective annihilation of the nonadiabatic coupling between the singlet and triplet states Gelin09 . It operates through the combined effect of the cavity field (through Rabi cycling) the system itself (through the intrinsic ${\rm S}_{1}^{(1)}-{\rm TT}^{(1)}$ CI), since the photonic mode does not directly couple the singlet and triplet states. Note also that the population evolution of the triplet states starts not from $t=0$, but from a characteristic time $\tau_{N}$ which increases with $N$. This indicates that the wavepacket needs $\approx\tau_{N}$ to reach the polaritonic CIs.
@@
Figure 7: Energy levels of two rubrene dimers in the cavity. The superscripts $(1)$ and $(2)$ number the dimers.
Let us now inspect evolutions $P_{\rm cav}(t)$ of the cavity-mode populations vs. the initial photon pumping numbers $N$ ($P_{\rm cav}(0)=N$) which deliver a complementary information. $P_{\rm cav}(t)$ for a single RD is depicted in Fig. 6(c). For $N$ from 1 to 6, the populations exhibit similar behavior: During the first 50 fs, they show several oscillations with a rough period of 22 fs which is related to the tuning mode. Then they reach a quasi-steady-state $\langle P_{\rm cav}\rangle$ value at longer times. It is worth mentioning that $\langle P_{\rm cav}\rangle$ are slightly larger than the initial populations $P_{\rm cav}(0)=N$. It reveals the population redistribution due to the coupling of the cavity mode and the ${\rm S}_{1}^{(1)}$ state. $P_{\rm cav}(t)$ for $N=7$ shows high-amplitude oscillations during the first $\sim 40$ fs, which matches the time $\tau_{7}$ during which the corresponding triplet population in Fig. 6(a) is almost zero. For $N=8$, the high-amplitude oscillations persist over the entire evolution of $P_{\rm cav}(t)$, which is a signature of the localization of the population on the photonic mode and the singlet state, which results in the SF stopping. Interestingly, the period of these fast oscillations reveals the effective Rabi frequency $\Omega_{N}\approx\Omega\sqrt{N}$ ($2\pi/\Omega_{8}\approx 7.3$ fs), which is defined under the assumption that the effective coupling to the cavity scales $\sim\sqrt{N}$.
Fig. 6(d) is a counterpart of Fig. 6(c), but for two RDs coupled to the cavity. For $N$ from 1 to 6, $P_{\rm cav}(t)$ reach their first local minima at $t\approx 17$ fs (dashed line in panel d), which indicates that a fraction of the photonic-mode population is transferred to the doubly excited electronic states of the RDs. Then, $P_{\rm cav}(t)$ exhibit oscillations with a period of $\sim 26$ fs, which cannot be directly related to the model parameters. In general, excitation to $1\leq N\leq 6$ reflects population transfer between photonic mode and excited states of the RDs. For $7\leq N\leq 9$, $P_{\rm cav}(t)$ shows fast Rabi oscillations at the initial stage, which correlate with the absence of the triplet excitation at the same time interval (cf. Fig. 6(b)). For $N=10$, fast Rabi oscillations spread over the entire $P_{\rm cav}(t)$ evolution, which is the signature of the SF switching-off in Fig. 6(b). In the long-time limit, the $P_{\rm cav}(t)$ reach quasi-steady states $\langle P_{\rm cav}\rangle$ the values of which are lower than the initial values $P_{\rm cav}(0)$. This indicates a net population transfer from the photonic mode to the doubly-excited electronic states of RDs.
@@
The above analyses demonstrate that the SF efficiency is primarily determined by two factors. The first factor is the position of the polaritonic CIs (see Figure 5). If the CIs are located in the Franck-Condon region, the subsequent population transfer is facilitated and the higher SF efficiency is achieved. The second factor is the initial wavepacket preparation and the state of the cavity which is controlled by the photonic-mode pumping number $|\mu_{1}(0)|^{2}=N$. This later parameter plays the crucial role, increasing the SF efficiency for small $N$ and fully suppressing SF once $N$ reaches a certain critical value. Similar results are obtained in Ref. Vendrell1 . If the number of photons is higher than the number of the available excited states, the effective cavity-matter coupling may increase as well, leading to slower motion of the nuclear wave packet, which gets partially trapped in the region where the molecular electronic energy gap becomes resonant with the cavity.
Furthermore, collective effects, such as polaron decoupling and collective dark states, have significant impacts on the cooperative polariton dynamics Gu2 . SF dynamical pathways can also be modified by strong photon-matter coupling. For instance, the SF channel via the UP state impedes vibronic losses to the dark-states manifold, which evidently enhances the SF efficiency Vendrell3 .
III.3 Spectroscopic signatures of Polaritonic CIs
@@
As aforementioned, strong coupling of RDs to the cavity mode drastically alters the SF efficiency owing to the cohesive action of the cavity-induced and intrinsic CIs. This processes are imprinted into nonlinear femtosecond spectroscopic responses of these systems. To find spectroscopic signatures of the polaritonic SF occurring in multidimensional PES landscapes, the signals must be accurately simulated and adequately interpreted. For this to achieve, it is essential to realistically model RDs. Indeed, all third-order spectroscopic responses consist of three contributions, the GSB, the SE, and the ESA Darius04 ; Leonas . The GSB and SE reflect the photoinduced processes in the manifolds with $N=0$ (the ground state) and $N=1$ (singly-excited states), while ESA involves also the doubly-excited states, $N=2$. Hence modeling individual RDs as systems containing only singly-excited states ${\rm S}_{1}^{(j)}$ and ${\rm TT}^{(j)}$ is insufficient: we have to include higher-excited states ${\rm S}_{n}^{(j)}$ and ${\rm TT}_{n}^{(j)}$ sun2 . Following Miyata ; sun2 , we set $E_{{\rm S}_{n}}^{(j)}=4.33$ eV and $E_{{\rm TT}_{n}}^{(j)}=4.68$ eV in our simulations. For two RDs in the cavity, we thus obtain a single ground state, 5 singly-excited states and 15 doubly-excited states which are specified in Fig. 7. The states, in turn, can be subdivided into the matter states ($n_{c}=0$, left part of Fig. 7) and hybrid states ($n_{c}>0$, right part of Fig. 7).